OpenStudy (anonymous):

What number is 5/6 of 2 2/5?

OpenStudy (anonymous):

first convert your fractions into decimals http://www.calculatorsoup.com/calculators/math/mixed-number-to-decimal-calculator.php

OpenStudy (jdoe0001):

\(\bf \cfrac{5}{6}\ of \ 2\frac{2}{5}\implies \cfrac{5}{6}\cdot 2\frac{2}{5}\)

OpenStudy (anonymous):

Then, use this http://www.percentagecalculator.net/